If the sum of first p terms of an A.P. is equal to the sum of the first q terms, then find the sum of the first (p + q) terms

Grade:11

5 Answers

vinay AskiitiansExpert-IITD
13 years ago

Dear Chanchal,

equate sum of first p and q terms then write d multiplied by its coefficients on one side and a and its coefficients on the other and cancel (p-q) factor, collect all terms one side and you will get expression similar to that of  p+q term's sum which will give you value of sum as 0. 

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.

rajan jha
13 years ago

A/Q,

  p/2(2a+(p-1)d)=q/2(2a+(q-1)d)

=)p(2a+(p-1)d)=q(2a+(q-1)d)

=)2ap+p(p-1)d=2aq+q(q-1)d

=)2ap-2aq=d(q(q-1)-p(p-1))

 =)2a(p-q)=d(q^2-p^2+p-q)

after solving you get,  2a=-(p+q-1)d............................(1)

 

NOW

sum upto (p+q) terms = (p+q)/2(2a+(p+q-1)d)

                                 =(p+q)/2(2a-2a)                                  (from 1)

                                 =0.

best of luck.........................

like my answer?

 

 

vikas askiitian expert
13 years ago

let the first term and common difference of AP is a and d....

sum of first q terms =sum of first p terms =k

q/2{2a+(q-1)d} = p/2{2a+(p-1)d} = k 

2a = 2k(q+p-1)/pq   and d=-2k/pq

now sum of first p+q terms =Sp+q= p+q/2 {2a +(p+q-1)d}

                                     =(p+q)/2{2k(p+q-1)/pq  - 2k(p+q-1)/pq }

                                    =0

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

one question to all of you solved this wat will be the sum when p-q=0 that is p=q this is the difficult case. why you neglected this case.

mycroft holmes
13 years ago

The sum to n terms of an AP is given by the general expression 

 

We have 

 

Since p,q are assumed unequal, we have  so that the sum of (p+q)

terms is 0